Job Cost Journal Entries and T Accounts

Following are certain operating data for Durango Manufacturing Company for January 2019:

Materials Inventory Work in Process Inventory Finished Goods Inventory
Beginning inventory $57,000 $24,000 $75,000
Ending inventory 33,000 40,500 48,000

Total sales were $1,800,000, on which the company earned a 40% gross profit. Durango uses a predetermined manufacturing overhead rate of 120% of direct labor costs. Manufacturing overhead applied was $360,000. Exclusive of indirect material used, total manufacturing overhead incurred was $243,000; it was over-applied by $22,500.

Required

Compute the following items. (Set up T accounts for Materials Inventory, Work in Process Inventory, Finished Goods Inventory, and Manufacturing Overhead; fill in the known amounts; and then use the normal relationships among the various accounts to compute the unknown amounts.)

Materials Inventory
Beg. bal.

57,000

Answer (d)
(f) Answer Answer (e)
End. bal. Answer
Work in Process Inventory
Beg. bal.

24,000

Answer (b)
(c) Answer
(d) Answer
Answer
End. bal. Answer
Manufacturing Overhead
Beg. bal.

0

Answer
Answer
(e) Answer
Answer End. bal.
Finished Goods Inventory
Beg. bal.

75,000

Answer (a)
(b) Answer
End. bal. Answer
Answer #1
Materials Inventory
Beg. bal.

57000

409500 (d)
(f) 480000 94500 (e)
End. bal. 33000
Work in Process Inventory
Beg. bal.

24000

1053000 (b)
(c) 300000
(d) 409500
360000
End. bal. 40500
Manufacturing Overhead
Beg. bal.

0

360000
243000
(e) 94500
22500 End. bal.
Finished Goods Inventory
Beg. bal.

75000

1080000 (a)
(b) 1053000
End. bal. 48000

As the ending balance of materials inventory, work in process inventory and finished goods inventory is given, we will fill them. Then the manufacturing overhead applied and actual manufacturing overhead except indirect materials is given we will also fill them.

(a) Cost of goods sold= Sales-Gross profit

= $1800000-(1800000*40%)= $1080000

(b) Cost of goods manufactured= Ending balance+Cost of goods sold- Beginning balance

= $48000+1080000-75000= $1053000

(c) Direct labor= Manufacturing overhead applied/120%

= $360000/120%= $300000

(d) Direct materials= Cost of goods manufactured+Ending balance-Beginning balance-Direct labor-Manufacturing overhead applied

= $1053000+40500-24000-300000-360000= $409500

(e) Indirect materials= Manufacturing overhead applied-Actual manufacturing costs-Overapplied overhead

= $360000-243000-22500= $94500

(f) Raw materials purchased= Direct materials+Indirect materials+Ending balance-Beginning balance

= $409500+94500+33000-57000= $480000